OpenStudy (anonymous):

Rationalize the denominator of radical 18 over radical 3?

OpenStudy (anonymous):

.. √18 / √3 = [ √18 * √3 ] / [ √3 * √3 ] = [ √(18 * 3) ] / 3 = [ √(6 * 3 * 3) ] / 3 = [ 3 √6 ] / 3 = √6

OpenStudy (anonymous):

can you rationalize 3 over radical 2?

OpenStudy (anonymous):

radical 3 two's cancel ( as long as the last two is not under the radical)

OpenStudy (anonymous):

\[\Large \frac32 \sqrt2\] this is correct.

OpenStudy (anonymous):

\[\frac{\sqrt{18}}{\sqrt{3}}=\sqrt{\frac{18}{3}}=\sqrt{6}\]no need for rationalizng denominator

OpenStudy (anonymous):

\[\frac{3}{\sqrt{2}}=\frac{3}{\sqrt{2}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{3\sqrt{2}}{2}\]this one does require rationalizing denominator

OpenStudy (anonymous):

3 \cdot2 = \[3 \cdot 2\] on latex ;) @satellite73